OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x+3y=5 and passes through the point (9,-3)

OpenStudy (amistre64):

parallel lines differ at most by a constant

OpenStudy (amistre64):

2x+3y=5 is parallel to 2x+3y=k input the given (x,y) values to determine the value of k

OpenStudy (anonymous):

I'm still lost >.<

OpenStudy (amistre64):

how?

OpenStudy (anonymous):

Cause I can't figure out the answer

OpenStudy (anonymous):

K = 5 I understand that much of what you wrote

OpenStudy (amistre64):

not quite ... K is some constant that needs to be determined by the point they give you. parallel lines SHARE much of the same equation, so they look almost exactly alike except of that K value

OpenStudy (anonymous):

Oh ok so how do I figure out the answer

OpenStudy (amistre64):

2x + 3y = 5 is parallel to 2x + 3y = K ; use the point (x,y) they give you to define the K do you know how to read a point?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then use the x and y value from the given point and work the multiplication and addition

OpenStudy (anonymous):

Yea now I'm lost again can you please just answer my question its the last question I need

OpenStudy (amistre64):

2x + 3y = K ; the given point is (9,-5) 2(9) + 3(-5) = K 18 - 15 = K

OpenStudy (anonymous):

So 3= k

OpenStudy (amistre64):

yes, which allows us to finish the line equation as: 2x + 3y = 3

OpenStudy (anonymous):

Actually it's a -3 not a -5 for the equation

OpenStudy (anonymous):

So technically it's 9